OpenStudy (shaniehh):

What is the value of the leading coefficient a if the polynomial function P(x) = a(x + b)2(x − c) has multiplicity of 2 at the point (–3, 0) and also passes through the points (2, 0) and (0, 36)?

OpenStudy (shaniehh):

–2 –3 3 36

OpenStudy (shaniehh):

@dan815 @zepdrix

OpenStudy (shaniehh):

@Whitemonsterbunny17 @pooja195 @triciaal @jigglypuff314 @Jaynator495

OpenStudy (shaniehh):

@Directrix

OpenStudy (campbell_st):

so the polynomial is \[P(x) = a(x + b)^2(x-c)\] is that correct...?

OpenStudy (campbell_st):

so the polynomial has factors \[(x + b)^2 \] this has multiplicity or 2 and has a zero at x + b = 0 or x = -b then linear factor is x - c = 0 so a zero exists at x = c compare this to the given information multiplicity 2 at (-3, 0) so x =-3 compare this to the zero above with multiplicty 2 x = -b then you can say b = 3 so the same for x = c you a told of a zero at (2, 0) or x = 2 so c = 2 so the polynomial is \[P(x) = a(x+ 3)^2(x - 2)\] now to find a, use the point (0,36) substitute x = 0 and P(x) = 36 then solvee for a hope it helps

OpenStudy (shaniehh):

so the answer is D @campbell_st